OpenStudy (anonymous):

what is an equation for the function. y=4/x

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

multiply both sides by x then subtract 4.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

right side: y x left side: \(\dfrac{4}{x}x\) = 4 put the 2 together y x = 4

OpenStudy (anonymous):

well the answer is y=\[y =\frac{ x-5 }{ x^2+4x+3}\] but i dont no how to get it

OpenStudy (anonymous):

er im sorry wrong one the answer is y=4/x+6 +2

OpenStudy (anonymous):

Then you wrote your question wrong. Check it please.

OpenStudy (anonymous):

@ DanielM_113